An organic compound contains 40.687% carbon, 5.085% hydrogen and rest is Oxygen. Calculate the molecular formula of the compound with molar mass=118
1
Expert's answer
2020-06-29T07:47:48-0400

Given % Of Carbon = 40.67 ;

% Of Hydrogen = 5.085 ;

% Of Oxygen = 54.245 ;


% mole Of Carbon = 40.67 /12 = 3.390 .

% mole Of Hydrogen = 5.085 /1 = 5.085 .

% mole Of Oxygen = 54.245 /16 = 3.390 .


Molar ratio Of Carbon : Hydrogen : Oxygen . = 3.390/3.390 : 5.085 /3.390 : 3.390 /3.390 :

Molar ratio Of Carbon : Hydrogen : Oxygen . = 1 : 1.5 : 1 .

Integral Molar Ratio = 2 : 3 : 2 .

So Empirical Formula = C2H3O2 ; ANSWER.